OpenStudy (anonymous):

solve for y: 5x - 15y = 20

OpenStudy (anonymous):

subtract 5x to both sides.

OpenStudy (alexandervonhumboldt2):

-15y=20-5x

OpenStudy (alexandervonhumboldt2):

y=-20/15+1/3x

OpenStudy (alexandervonhumboldt2):

y=-4/3+1/3x

OpenStudy (anonymous):

yep then divide both sides by -15

OpenStudy (anonymous):

so the answer is?

OpenStudy (alexandervonhumboldt2):

y=1/3x-4/3